OpenStudy (anonymous):

The theoretical probability of undesirable side effects resulting from taking Grebex is 1 in 10. 100 people take Grebex to lower their blood pressure, how many will encounter undesirable side effects?

OpenStudy (helder_edwin):

u have a probability of 1 in 10 that is 1/10 what do u think u have to do about the 100?

OpenStudy (anonymous):

1/10=x/100---> x=10 people would encouter side effects

OpenStudy (helder_edwin):

yes also the number of people that would ..... is \(1/10\cdot100=10\)

OpenStudy (anonymous):

thanks

OpenStudy (helder_edwin):

u r welcome